Fisk Corporation is trying to improve its inventory control system and has installed an online computer at its retail stores. Fisk anticipates sales of 84,500 units per year, an ordering cost of $12 per order, and carrying costs of $1.20 per unit.


 a. What is the economic ordering quantity? 

    Economic ordering quantity = _______ units 

 b. How many orders will be placed during the year? 

    orders Number of = _______ orders 

 c. What will the average inventory be? 

    Average inventory = _______ units 

 d. What is the total cost of ordering and carrying inventory? 

    Total cost = _______ 
Answer #1

a)

Economic order quantity

= Sqrt (2 x A x O / C)

Where,

A = Annual Demand = 84,500

O = Ordering cost per order = $12

C = Carrying cost per unit per annum = $1.20

So, EOQ

= Sqrt(2 x 84,500 x 12 / 1.20)

= Sqrt(1,690,000)

= 1,300

b)

Number of orders

= Annual Demand / EOQ

= 84,500 / 1,300

= 65 orders

c)

Average Inventory

= EOQ / 2

= 1,300 / 2

= 650

d)

Total cost of ordering and carrying

Ordering cost = Number of orders x Ordering cost per order

= 65 x $12

= $780

Carrying cost

= Average inventory x Carrying cost per unit per annum

= 650 x $1.20

= $780

So, Total cost of ordering and carrying

= $780 + $780

= $1,560